Two firms Intel and AMD in the CPU market have combined demand given by Q = 100 - P. Their total costs are given by TC Intel = 4 Q Intel + 2 Q Intel^2 and TC AMD = 4 Q AMD + 2 Q AMD^2. If they successfully collude, the market price will be??
Answer #1

2 992 = 100-P 8 = Sintel & SAMO Tointel = YQIT 281 TCAMD = USA + 282 To = 48+ 28² 4+49 MC = MR = 100-29 MC = = MR 4+48= 100-2